OpenStudy (anonymous):

The function H(x) is described below. H(x) = 3x - 4 What is the value of H(6)? A. 5 B. 6 C. 14

OpenStudy (anonymous):

H(x) = 3x - 4 H(6) = 3 * 6 -4

OpenStudy (anonymous):

Just Replace x with 6

OpenStudy (anonymous):

H(6) = 18 - 4 = 14